# Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic

### v1

@@ 6 Discussion
Grokking, Benign Overfitting, and Implicit Bias.
Recently, [94] connects the grokking phenomenon to benign overfitting [7, 11, 84, 24, 26]. It shows how the network undergoes a grokking period from catastrophic to benign overfitting. [47] uses implicit bias [72, 28, 41, 79, 59, 10, 51, 38, 92, 70, 74, 93] to explain grokking, where grokking happens if the early phase bias implies an overfitting solution while late phase bias implies a generalizable solution. The intuition from the benign overfitting and the implicit bias well align with our observation in Section 5.3. It is interesting and valuable to rigorously analyze the grokking or emergent ability under different function class complexities, e.g., Eq (1). We leave this challenge problem as a future work.
@@ Acknowledgement
In Appendix A, we introduce some definitions that will be used in the proof. In Appendix B, we introduce some auxiliary lemma from previous work that we need. In Appendix C, Appendix D, Appendix E, Appendix F, we provide the proof of our Lemmas and our main results. In particular, we provide two versions of proof (1) $k=3$ and (2) general $k\geq 3$. We use $k=3$ version to illustrate our proof intuition and then extend our proof to the general $k$ version.
Appendix A Notations and Definitions
@@
We denote $\mathbb{Z}_{p}$ as the modular group on $p$ integers, e.g., $\mathbb{Z}_{p}=[p]$, where $p>2$ is a given prime number. We denote $\mathcal{X}:=\mathbb{Z}_{p}^{k}$ as the input space and $\mathcal{Y}:=\mathbb{Z}_{p}$ as the output space. We denote $D_{p}:=\{((a_{1},\dots,a_{k}),\sum_{i\in[k]}a_{i}):a_{1},\dots,a_{k}\in\mathbb{Z}_{p}\}$ as the modular dataset.
Suppose we have a $L_{2,k+1}$ (matrix) norm $\|\cdot\|_{2,k+1}$ (Definition 3.3) and a class of parameterized functions $\{f(\theta,\cdot)\mid\theta\in\mathbb{R}^{U}\}$, where $f:\mathbb{R}^{U}\times\mathcal{X}\rightarrow\mathbb{R}^{\mathcal{Y}}$ and a matrix set $\Theta:=\{\|\theta\|_{2,k+1}\leq 1\}$. We consider single-hidden layer neural networks with polynomial activation functions and without biases. With parameters $\theta\in\Theta$, we denote $f(\theta,x)$ as the network’s output for an input $x$. For one-hidden layer networks, we denote $f$ as:
$f(\theta,x):=\sum_{i=1}^{m}\phi(\theta_{i},x),$
where $\theta:=\{\theta_{1},\ldots,\theta_{m}\}$. Here, we denote $\phi$ as one neuron and $\theta_{i}\in\Omega$ are the corresponding weights in that neuron, where $\Omega$ is a vector set. Let $\Omega^{\prime}$ be a subset of $\Omega$. For every $i\in[m]$, when either there exists $\alpha_{i}>0$ such that $\alpha_{i}\theta_{i}\in\Omega^{\prime}$ or $\theta_{i}=0$, then we say the parameter set $\theta=\{\theta_{1},\ldots,\theta_{m}\}$ has directional support on $\Omega^{\prime}$. A single neuron is represented as:
$\phi(\{u_{1},\dots,u_{k},w\},x_{1},\dots,x_{k}):=~{}(u_{1}^{\top}x_{1}+\dots+u_{k}^{\top}x_{k})^{k}w,$
where $u_{1},\dots,u_{k},w\in\mathbb{R}^{p}$ are the neuron’s weights, and $x_{1},\dots,x_{k}\in\mathbb{R}^{p}$ are the network inputs. Inputs $x_{1},\dots,x_{k}$ are one-hot vectors representing group elements. For input elements $(a_{1},\dots,a_{k})$, a neuron simplifies to
$\phi(\{u_{1},\dots,u_{k},w\},a_{1},\dots,a_{k})=(u_{1}(a_{1})+\dots+u_{k}(a_{k}))^{k}w,$
with $u_{1}(a_{1})$ being the $a_{1}$-th component of $u_{1}$. The output of the neuron is in $p$-dimension for cross-entropy loss. With $\theta=\{u_{i,1},\dots,u_{i,k},w_{i}\}_{i=1}^{m},$ the network is denoted as:
$~{}f(\theta,a_{1},\dots,a_{k}):=~{}\sum_{i=1}^{m}\phi(\{u_{i,1},\dots,u_{i,k},w_{i}\},a_{1},\dots.a_{k}).$
@@
Definition A.2.
We define $\Theta^{*}$ as follows
  • $\Theta^{*}:=\arg\max_{\theta\in\Theta}h(\theta)$.
Appendix B Tools from Previous Work
Section B.1 states that we can use the single neuron level optimization to get the maximum-margin network. Section B.2 introduces the maximum-margin for multi-class.
B.1 Tools from Previous Work: Implying Single/Combined Neurons
Lemma B.1 (Lemma 5 in page 8 in [55]).
@@ B.1 Tools from Previous Work: Implying Single/Combined Neurons
B.2 Tools from Previous Work: Maximum Margin for Multi-Class
Lemma B.2 (Lemma 6 in page 8 in [55]).
@@ B.2 Tools from Previous Work: Maximum Margin for Multi-Class
  • Suppose that $\exists\{\theta^{*},q^{*}\}$ such that Equations (4) and (11), and 3.10 holds.
@@
Appendix C Class-weighted Max-margin Solution of Single Neuron
Section C.1 introduces some definitions. Section C.2 shows how we transfer the problem to discrete Fourier space. Section C.3 proposes the weighted margin of the single neuron. Section C.4 shows how we transfer the problem to discrete Fourier space for general $k$ version. Section C.5 provides the solution set for general $k$ version and the maximum weighted margin for a single neuron.
C.1 Definitions
Definition C.1.
@@ C.1 Definitions
Definition C.2.
When $k=3$, provided the following conditions are met
@@
C.2 Transfer to Discrete Fourier Space
@@
  • We define $*\Omega_{q}^{{}^{\prime}*}$ in Definition C.2.
@@ C.3 Get Solution Set
  • We define $*\Omega_{q}^{{}^{\prime}*}$ in Definition C.2.
@@
By Lemma C.3, we only need to maximize Equation (C.2).
@@
Consider the odd $p$, Equation (C.2) becomes:
@@
where the first step comes from definition (C.2), the second step follows from Eq. (14), the third step comes from $\widehat{u}_{1}(-j)=\overline{\widehat{u}_{1}(j)}$ and $\widehat{u}_{1}(j)=|\widehat{u}_{1}(j)|\exp(\mathbf{i}\theta_{u_{1}}(j))$, the last step follow from Euler’s formula.
@@
C.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition C.5.
@@ C.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition C.6.
Provided the following conditions are met
@@
  • We define $*\Omega_{q}^{{}^{\prime}*}$ in Definition C.6.
@@
C.5 Get Solution Set for General $k$ Version
Lemma C.8 (Formal version of Lemma 4.2).
@@ C.5 Get Solution Set for General $k$ Version
  • Let $*\Omega_{q}^{{}^{\prime}*}$ be defined as Definition C.6.
@@
By Lemma C.7, we only need to maximize Equation (C.4). Thus, the mass of $\widehat{u}_{1},\dots,\widehat{u}_{k}$, and $\widehat{w}$ must be concentrated on the same frequencies. For all $j\in\mathbb{Z}_{p}$, we have
@@
Considering odd $p$, Equation (C.4) becomes:
@@
where the first step follows from definition (C.4), the second step comes from Eq. (23), the last step follows from Eq. (24), i.e., Euler’s formula.
@@
Appendix D Construct Max Margin Solution
Section D.1 proposed the sum-to-product identities for $k$ inputs. Section D.2 shows how we construct $\theta^{*}$ when $k=3$. Section D.3 gives the constructions for $\theta^{*}$ for general $k$ version.
D.1 Sum-to-product Identities
Lemma D.1 (Sum-to-product Identities).

### v2

@@ 6 Discussion
Grokking in Large Language Models.
The interpretability of grokking in large language models (LLMs) is explored in [70]. By examining various intermediate states within the residual stream of the Transformer model, it is validated that the model employs Fourier features to tackle the modular addition task. However, fully comprehending how the Transformer model and LLMs perform modular addition remains challenging based on the current work, particularly from a theoretical standpoint. We contend that beginning with a simplistic model setup and achieving a thorough and theoretical understanding of how the network utilizes Fourier features to address the problem serves as a valuable starting point and it provides a theoretical understanding of the grokking phenomenon. We believe that further study on LLMs will be an interesting and important future direction.
Grokking, Benign Overfitting, and Implicit Bias.
Recently, [115] connects the grokking phenomenon to benign overfitting [9, 13, 99, 26, 28]. It shows how the network undergoes a grokking period from catastrophic to benign overfitting. [55, 52] uses implicit bias [85, 32, 48, 92, 69, 12, 59, 44, 113, 114] to explain grokking, where grokking happens if the early phase bias implies an overfitting solution while late phase bias implies a generalizable solution. The intuition from the benign overfitting and the implicit bias well align with our observation in Section 5.3. It is interesting and valuable to rigorously analyze the grokking or emergent ability under different function class complexities, e.g., Eq (1). We leave this challenge problem as a future work.
@@ Acknowledgement
Roadmap.
In Section A, we provide the potential limitations of this work. In Section B, we discuss the societal impacts of our work. In Section C, we introduce some definitions that will be used in the proof. In Section D, we introduce some auxiliary lemma from previous work that we need. In Section E, Section F, Section G, Section H, we provide the proof of our Lemmas and our main results. In particular, we provide two versions of proof (1) $k=3$ and (2) general $k\geq 3$. We use $k=3$ version to illustrate our proof intuition and then extend our proof to the general $k$ version. Finally, in Section I, we provide more experimental results and implementation details.
Appendix A Limitations
Our work has made progress in exploring how neural networks and Transformers can solve complex mathematical problems such as modular addition operation, but the practical application scope of their conclusions is limited.
On the other hand, we admit that our theorem can provide intuition but cannot fully explain the phenomena shown in Figure 4. Thus, we would like to introduce this more general data setting to the community so that we can study and understand grokking in a more broad way. Studying the relationship between the number of neurons and the grokking strength is interesting and important, and we will leave it as our future work.
Appendix B Societal Impact
Our work aims to understand the potential of large language models in mathematical reasoning and modular arithmetic. Our paper is purely theoretical and empirical in nature (mathematics problem) and thus we foresee no immediate negative ethical impact.
We propose that neural networks and transformers prefer to learn Fourier circuits when training on modular addition involving $k$ inputs under SGD, which may have a positive impact on the machine learning community. We hope our work will inspire effective algorithm design and promote a better understanding of large language models learning mechanisms.
Appendix C More Notations and Definitions
@@
We define the vector norm and matrix norm as the following.
Definition C.1 ($L_{b}$ (vector) norm).
Definition C.2 ($L_{a,b}$ (matrix) norm).
@@
Definition C.4.
We define $\Theta^{*}:=\arg\max_{\theta\in\Theta}h(\theta)$.
Finally, let $\Omega:=\mathbb{R}^{p\times(k+1)}$ denote the domain of each $\theta_{i}$, and let $\Omega^{\prime}$ be a subset of $\Omega$. We say the parameter set $\theta=\{\theta_{1},\ldots,\theta_{m}\}$ has directional support on $\Omega^{\prime}$, if for every $i\in[m]$, either $\theta_{i}=0$ or there exists $\alpha_{i}>0$ such that $\alpha_{i}\theta_{i}\in\Omega^{\prime}$.
Appendix D Tools from Previous Work
Section D.1 states that we can use the single neuron level optimization to get the maximum-margin network. Section D.2 introduces the maximum-margin for multi-class.
D.1 Tools from Previous Work: Implying Single/Combined Neurons
Lemma D.1 (Lemma 5 in page 8 in [64]).
@@ D.1 Tools from Previous Work: Implying Single/Combined Neurons
D.2 Tools from Previous Work: Maximum Margin for Multi-Class
Lemma D.2 (Lemma 6 in page 8 in [64]).
@@ D.2 Tools from Previous Work: Maximum Margin for Multi-Class
  • Suppose that $\exists\{\theta^{*},q^{*}\}$ such that Equations (4) and (5), and 3.8 holds.
@@
Appendix E Class-weighted Max-margin Solution of Single Neuron
Section E.1 introduces some definitions. Section E.2 shows how we transfer the problem to discrete Fourier space. Section E.3 proposes the weighted margin of the single neuron. Section E.4 shows how we transfer the problem to discrete Fourier space for general $k$ version. Section E.5 provides the solution set for general $k$ version and the maximum weighted margin for a single neuron.
E.1 Definitions
Definition E.1.
@@ E.1 Definitions
Definition E.2.
Let $\eta$ be defined in Definition E.1. When $k=3$, provided the following conditions are met
@@
E.2 Transfer to Discrete Fourier Space
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@ E.3 Get Solution Set
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@
By Lemma E.3, we only need to maximize Equation (E.2).
@@
Consider the odd $p$, Equation (E.2) becomes:
@@
where the first step comes from definition (E.2), the second step follows from Eq. (8), the third step comes from $\widehat{u}_{1}(-j)=\overline{\widehat{u}_{1}(j)}$ and $\widehat{u}_{1}(j)=|\widehat{u}_{1}(j)|\exp(\mathbf{i}\theta_{u_{1}}(j))$, the last step follow from Euler’s formula.
@@
E.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition E.5.
@@ E.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition E.6.
Let $\eta$ be defined in Definition E.5. Provided the following conditions are met
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.6.
@@
E.5 Get Solution Set for General $k$ Version
Lemma E.8 (Formal version of Lemma 4.2).
@@ E.5 Get Solution Set for General $k$ Version
  • Let $\Omega_{q}^{{}^{\prime}*}$ be defined as Definition E.6.
@@
By Lemma E.7, we only need to maximize Equation (E.4). Thus, the mass of $\widehat{u}_{1},\dots,\widehat{u}_{k}$, and $\widehat{w}$ must be concentrated on the same frequencies. For all $j\in\mathbb{Z}_{p}$, we have
@@
Considering odd $p$, Equation (E.4) becomes:
@@
where the first step follows from definition (E.4), the second step comes from Eq. (17), the last step follows from Eq. (18), i.e., Euler’s formula.
@@
Appendix F Construct Max Margin Solution
Section F.1 proposed the sum-to-product identities for $k$ inputs. Section F.2 shows how we construct $\theta^{*}$ when $k=3$. Section F.3 gives the constructions for $\theta^{*}$ for general $k$ version.
F.1 Sum-to-product Identities
Lemma F.1 (Sum-to-product Identities).